HMRC report that 5% of taxpayers filling out the short form make serious mistakes. In a survey, 50 of these forms are chosen at random.

Find the probability that at most two forms have serious mistakes. Perform the calculation:

  1. using the Binomial distribution,
  2. using the Poisson approximation to the binomial distribution.

Discuss whether you consider the approximation is appropriate.
Answer #1

a)

Let X denote the number of forms that contain mistakes. Then X ~ Bin(n = 50, p=0.05)

P(X = ) = (0.05)(1 – 0.05)50-1, 1 = 0, 1, 2, ...., 50

Required probability = P(X < 2) = P(X =0) + P(X = 1) + P(X = 2)

(0.05)º(1-0.05)50-0 + (0,0511(1_0 05:50-1 (0.05)(1– (1) 0.05)50-2

= 0.5405

b)

Using Poisson approximation, X ~ Poi(X= np = 50(0.05) = 2.5)

2.54e-2.5 P(X = 1) = — ,I=0,1,2,3,..

Required probability = P(X < 2) = P(X =0) + P(X = 1) + P(X = 2)

=\frac{2.5^0e^{-2.5}}{0!}+\frac{2.5^1e^{-2.5}}{1!}+\frac{2.5^2e^{-2.5}}{2!}=0.5438

If 'n' is greater than 20 and 'np' is less than 5, then Poisson approximation is a good approximation.

Here, n=50 and np=2.5. Hence, the approximation is appropriate.